OpenStudy (anonymous):

Janine made a cylindrical vase in which the sum of the lateral area and area of one base was about 3000 square centimeters. The vase had a height of 50 centimeters. Find the radius of the vase. Explain the method you would use to find the radius.

OpenStudy (anonymous):

He called you out outkast

OpenStudy (anonymous):

who called me out?

OpenStudy (anonymous):

what are u talking about

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

direct please be my fan! I follow back!

Directrix (directrix):

Are we working math problems here or having a fight? I am here for the math and if that is not happening, I'm moving on to another problem.

OpenStudy (anonymous):

is this right 3000=( 2*Pi*r*h)+Pi*r^2 3000= (2*Pi*r*50) +Pi*r^2 3000=100Pir+Pir^2 3000=Pir(100+r)

OpenStudy (anonymous):

Nah man

OpenStudy (anonymous):

but i don't know how to continue

Directrix (directrix):

Does the answer need to be exact or will an approximation do? It's quadratic formula to solve for r. @malkassir

OpenStudy (anonymous):

yeah thnx man

Directrix (directrix):

I think I missed an "r." You did great on the equation.

Directrix (directrix):

Here's the Wolf's approximations.

OpenStudy (anonymous):

this is so great thank u

Directrix (directrix):

I might stop here: 3000=100Pir+Pir^2 pi r^2 + 100 pi r - 3000 = 0 It appears that we will need to factor or use the quadratic formula to get r.

Directrix (directrix):

Glad to help. Check my calculation. You did the heavy work getting the equation written.

OpenStudy (anonymous):

can u help me on the other question

Directrix (directrix):

Where is the other question?